explain why there is no value of x for which the expressions x + 3 and x + 4 are equal?
My 7th grade son has this question on one of his worksheets and sadly I am unable to help him.
Thanks so much in advance for your help

Found 2 solutions by checkley71, Edwin McCravy:
Answer by checkley71(8403)

X+3=X+4
X-X=4-3
0=-1 SAYS THAT THERE IS NO SOLUTION TO THIS PROBLEM.
ALSO: NO NUMBER INCREASED BY 3 CAN BE EQUAL TO THE SAME NUMBER INCREASED BY 4!

Answer by Edwin McCravy(20056)

explain why there is no value of x for which the expressions x + 3 and x + 4
are equal?
My 7th grade son has this question on one of his worksheets and sadly I am
unable to help him.
Thanks so much in advance for your help
------------------------------

If there were such a value for x, then you could 
subtract x from both sides like this:

                  x + 3 = x + 4
                 -x      -x
              -------------------
                  0 + 3 = 0 + 4
                      3 = 4

That is, if there were such a value for x then 
3 would have to equal to 4. But 3 does not equal
4, so therefore assuming there was such a value 
for x is wrong.  Therefore there can be no such 
value for x.

Edwin